By how much must inventory change (in dollars) for the current ratio to equal 1?
Cash      3,000,000 Accounts Payable      4,300,000
Accounts Receivable      1,000,000 Notes Payable      5,100,000
Inventory      5,000,000

a. $0

b. $100,000

c. $3,400,000

d. $400,000
Answer #1
Existing current assets = 3000000+1000000+5000000 = $ 90,00,000
Existing current liabilities = 4300000+5100000 = $ 94,00,000
Required current assets for current ratio to be 1 = $ 94,00,000
Less: Existing current assets $ 90,00,000
Increase in inventory required $    4,00,000
Answer: Option [d] $400,000